OpenStudy (g33k):

FAN AND MEDAL FOR HELP PLEASE Tom recorded his lunch expenditure each day for one week in the table below. Find the mean, standard deviation, and variance of Tom's lunch expenditures. Round to the nearest thousandth. Day: Sunday | Monday | Tuesday | Wednesday | Thursday | Friday | Saturday Expenditure: $4.85 | $5.10 | $5.50 | $4.75 | $4.50 | $5.00 | $6.00

OpenStudy (mrnood):

Start with the mean Add all the numbers together and divide by the number of days Post your answer here and I will help with the next step..

OpenStudy (mrnood):

can you add toghetehr the 6 numbers? when you've done tha divide by 6 that is the value for the mean Post your answer here and I will help with the next step..

OpenStudy (g33k):

okay, hang on

OpenStudy (g33k):

okay would the mean be 5.1?

OpenStudy (mrnood):

No

OpenStudy (mrnood):

OOOOPS Sorry - my mistake Yes - the mean IS 5.1

OpenStudy (g33k):

i did, i got 5.1 $4.85 + $5.10 + $5.50 + $4.75 + $4.50 + $5.00 + $6.00 = 35.7 / 7 = 5.1

OpenStudy (g33k):

oh okay

OpenStudy (mrnood):

Now you need to work out the difference between each number and teh mean e.g. for ther first one: 4.85 - 5.1 = -.25 and then SQUARE that number (4.85-5.1)^2 = 0.0625 write a list of the 7 values you get from those calculations